OpenStudy (anonymous):

NEED DESPARATE HELP, WILL FAN AND MEDAL BEST ANSWER! Question 3.3. Which equation is equivalent to x + 3y - z = 3? (Points : 1) 3x + 6y - 3z = 6 4x + 12y - 4z = 12 2x + 4y - 2z = 8 2x + 3y - 4z = 12

Directrix (directrix):

Take the given equation and compare it with each option one by one

OpenStudy (anonymous):

@Directrix seems like the first or third one? I think option 3....

Directrix (directrix):

x + 3y - z = 3 3x + 6y - 3z = 6 Is there a number you can multiply each of the terms of the first equation by to get the second equation?

Directrix (directrix):

^^ is the first option so answer the question.

Directrix (directrix):

What times x = 3x ?

OpenStudy (anonymous):

3 @Directrix

Directrix (directrix):

Does 3 * 3y = 6y ?

OpenStudy (anonymous):

no.. so its not option 1 @Directrix

OpenStudy (anonymous):

oh wait its totally option 2 isnt it!! @Directrix

Directrix (directrix):

Correct. Move to option 2 x + 3y - z = 3 4x + 12y - 4z = 12 What times x = 4x?

OpenStudy (anonymous):

cause times 4, everything matches!

Directrix (directrix):

Correct.

OpenStudy (anonymous):

wow thats so easy now that i think about it haha, thanks!

Directrix (directrix):

So, they are equivalent equations.

OpenStudy (anonymous):

thanks!!:)

Directrix (directrix):

You are welcome.